OpenStudy (anonymous):

What is a quartic function with only the two real zeros given? x = -4 and x = -1 A. y = x4 + 5x3 + 5x2 + 5x + 4 B. y = x4 - 5x3 - 5x2 - 5x - 4 C. y = -x4 + 5x3 + 5x2 + 5x + 4 D. y = x4 + 5x3 + 5x2 + 5x - 5

OpenStudy (anonymous):

im thinking the answer is B

OpenStudy (lastdaywork):

substitute x=-1 in each option and tell me those with y=0

OpenStudy (anonymous):

Ummmmm *le stupid* Can you help walk me through one of the answers to find y=0 and i can do the same with the others...

OpenStudy (lastdaywork):

"B. y = x4 - 5x3 - 5x2 - 5x - 4" y = (-1)^4 - 5(-1)^3 - 5(-1)^2 - 5(-1) - 4 y = 1 + 5 - 5 + 5 - 4 y = 2 Barter system - "le stupid" Name the language and convert it into English XD XD XD

OpenStudy (anonymous):

Its just something I say XDD and thanks

OpenStudy (lastdaywork):

Not the answer I was looking for :'(

OpenStudy (anonymous):

XDD I'm writing the process and I'm going to use it for the other answers to find which one is correct